40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Jan 25 13:02:46 2021, aaron, Update, PSOMA,  
    Reply  Thu Feb 18 13:05:50 2021, aaron, Update, PSOMA,  beat_ASD.pdflasers_ASD.pdf
Message ID: 2612     Entry time: Mon Jan 25 13:02:46 2021     Reply to this: 2652
Author: aaron 
Type: Update 
Category: PSOMA 
Subject:  

enter Mon Jan 25 13:02:47 2021

cantilever Q materials

I moved the following materials to measure cantilever Qs from cryo lab to QIL. The are located in bins labelled 'cantilever Q' in a cabinet on the S wall of the QIL lab (see photos).

  • cantilevers
  • HeNe laser and power
  • QPD from the GeNS setup, and power
  • What I think are the clamps Zach was using for cantilever Q measurements.

I also identified in QIL the IR labs cryostat that was used for cantilever Qs, and is now set up for PD measurements. I tried lifting out the cold plate, but felt resistance while opening and stopped. Probably the electrical connections are keeping the cold plate in place, and I would like to check with the user of this cryostat before proceeding.

three corner hat

I am recording spectra of the beat note between three lasers: PSOMA north laser, PSOMA south laser, and cryo cavities East laser. I unplugged the E laser from its fiber-to-free launch (after Faraday isolator), and instead send it to the PSOMA table by fiber. Without changing settings on the cryo cavs TEC or laser driver, the E laser is 8.1 mW when it reaches PSOMA table. When measuring the beat note between the cryo cavs E laser and PSOMA north (south) laser, I send the E laser through the south (north) 90-10 pickoff BS in place of the PSOMA laser. Before recording spectra, I tune the current of the PSOMA N and S lasers until all beat notes are < 20 MHz (there is a few MHz drift during the measurements).

Summary of the measurements, all taken at the FC 1611 (AC measurements on the Moku):

  • acquisition speed: 125 kHz
  • phaselock bandwidth: 10 kHz
  • coupling: AC, 50 Ohm
  • Moku input range: 1 Vpp
  • Measurement time: 1 m 40 s
  • freewheeling: yes
Pair of lasers DC intensity mon (mW) beat note power (dBm) beat frequency (MHz)
PSOMA north and south 6.0 3 18
PSOMA north, cryo cavs east 4.3 -1 24
PSOMA south, cryo cavs east 3.5 -2 5

I notice that the FC 1611 has a UPC connector on its face... but the rest of our components have APC connectors. I'll watch out for effects from this, and order an adapting fiber. I also see sidebands at the harmonics of the beat note when the note is > 0 dBm (linear regime of 1611 is only 1mW).

The rate-limiting step is transferring data from Moku RAM to the ipad over the Wifi... should set up the Moku on ethernet, and check out ctn-scripts for automating the process.

exit Mon Jan 25 16:10:59 2021

ELOG V3.1.3-